The data shown below describe children in four countries in terms of adequate dental care. What is the probability that a child from the United States has adequate dental care? What is the probability that a child from France has adequate dental care? What proportion of children have inadequate dental care? What proportion of children with inadequate dental care are from the United States?

New Drug 200 14% Placebo 200 19% Sample size Progression to hypertensiorn
Answer #1
Adequate Dental Care Inadequate Dental Care
UK 2539 176 2715
US 4563 209 4772
France 1298 307 1605
Mexico 2874 256 3130
Total 11274 948 12222

a) probability that a child from the US has adequate dental care =4563/4772=0.9562

b)probability that a child from France had adequate dental care=1298/1605=0.8087

c) proportion of children have inadequate dental care =11274/12222=0.9224

d)proportion of children with inadequate dental care are from the US=209/948=0.2205